i dont understand how i can use the square root. when you answer my question if you can put a few exersise and explain how i can do that, maube i can understand better
thanks
michelle

Answer by mszlmb(115)

The square root of a number is a number, which, when multiplied by itself, equals that number. For example, the square root of 25 is 5. Why? Because 5*5=25. The square root of 36 is what? 6. Because 6*6=36.
OK now how do we find square roots?
Let's look at the number 4096. Too big for now, you say? ... fine...whatever i didn't even like that number :P
ok fine we'll start with 144 (whiney!) ok
you put the number 144 on the right
144. WE know that 144 is divisible by 2... how? Oh oops I forgot to show you the table! here it is. this table explains how you know if a number is evenly divisible by one of these numbers:
1: all numbers are divisible by 1
2: number must end with an even digit (12,14,199948576 would work, 12.4,3,1 won't)
3: sum of digits of number add up to something divisible by 3 (ie 33,342 sum of digits is 3+3+3+4+2=15 1+5=6 so 33,342 is divisible by 3)
4: if the TENS digit is odd (1,3,5,7,9), then the ONES digit must be 2 or 6. and if the TENS digit is even (0,2,4,6,8) the ONES digit must be 0,4,8. (ie: 1124 has 10's dgt 2, even and 1's dgt 4. according 2 the rules it is evenly divisible; the answer is 1124/4=881)
5: 1's dgt is 0 or 5.
6: 1's digit is even and sum is divisible by 3
7: find one!
8: if 10's digit is 0, 8, or 4; 1's digit must be 8. if 10's digit is 1 or 9, 1's dgt must b 6. if 10's dgt is 2 or 6, 1's dgt must b 4. if 10's dgt is 3 or 7, 1's dgt must equal 2. if 10's dgt is 5 or 8, 10's dgt must b 0. if 10's dgt is 9, 1's is 0 or 8. if there is an odd number in the 100's place, add 4 to all sums unless there is an odd number in the thousands etc. o jeez this is nuts haha i'll organize it sum other day
9: sum divisible by 9
10: ends in 0.
OK that was exasporating; now
take 144/2=72/2=36/2=18/2=9/3=3/3=1
ok so we have 2*2*2*2*3*3
cross out one of each, because we are square rooting, and we have 2*2*3. the answer: 12. yea
to know how to square root any number, like 2, extensive research must be done on the equation of sqare rooting. its hard but probably worth it or somthin
wen they didn't have calculators, they memorized many square roots!
ok so now, the big #, 4096.
4096/2=2048/2=1024/2=512/2=256/2=128/2=64/2=32/2=16/2=8/2=4/2=2/2=1
ok so we have 2*2*2*2*2*2*2*2*2*2*2*2 cut out half
2*2*2*2*2*2=4*4*4=16*4=64
c that wasn't so hard the square root of 4096 is 64.
we've only just scratched the ice of an iceberg, IM me if u need more help.